OpenStudy (rainbow_dash):

A student guesses the answers to 6 questions on a true-false quiz. Find the probability that the indicated number of guesses are correct: at least 3.

OpenStudy (anonymous):

Here: n=6 , p=1/2

OpenStudy (anonymous):

So, u need to find P(x>=3)

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

@Rainbow_Dash ???

OpenStudy (rainbow_dash):

okay sure

OpenStudy (rainbow_dash):

so p = .5, so \[.5 \ge 3\], 3/.5 =6

OpenStudy (anonymous):

P(x>=3)=P(x=3) + P(x=4)+P(x=6)